OpenStudy (anonymous):

8x^2(9x^3+4x^2-3x+6)

OpenStudy (anonymous):

72x^5 + 32x^4 - 24x^3 + 48x

OpenStudy (gw2011):

72x^5+32x^4-24x^3+48x^2

OpenStudy (anonymous):

(-5)^-3

OpenStudy (gw2011):

(-5)^-3=1/(-5)^3=1/(-5)(-5)(-5)=-1/125